OpenStudy (anonymous):

need help y=x+z, x=9, z=11 what is the value

OpenStudy (lgbasallote):

y = x + z x = 9 and z = 11 therefore... y = 9 + 11 does that help?

OpenStudy (anonymous):

s that the value

OpenStudy (anonymous):

so is 9+11 th value

OpenStudy (lgbasallote):

what's 9+11

OpenStudy (anonymous):

20 is the value

OpenStudy (lgbasallote):

there you go

OpenStudy (anonymous):

oh thanks so much